# Theoretical Foundations of Prompt Engineering: From Heuristics to Expressivity

### v1

@@ 2.1 Universal Approximation and Expressivity Foundations
A standard entry point to expressivity is universal approximation: a model class is universal if its realizable functions are dense in a target space such as $C(K)$ for compact $K$. Recent theory refines this classical statement in directions that are directly relevant to prompt-based switching. One line of work replaces qualitative density with quantitative approximation guarantees, relating the error to architectural capacity (depth/width) and to function-space regularity (e.g., Besov-type smoothness classes for ReLU networks) (siegel2023optimal). Another line identifies minimality conditions and tight thresholds for universality, clarifying when universality fails due to insufficient width even on compact domains (kim2023minimum). Universality has also been studied beyond the compact setting via uniform approximation frameworks on broader domains, where noncompactness and tail control become essential (bournez2025universal), and via alternative parameterizations that re-express approximation mechanisms while retaining universal approximation goals (liu2024kan).
A complementary expansion of the UAT viewpoint concerns operator learning, where the target is a map between function spaces rather than $\mathbb{R}^{d}\to\mathbb{R}$. Neural-operator theory elevates universality to the operator level and connects approximation theory to PDE solution operators (kovachki2023neural), with recent results emphasizing structural properties (e.g., injectivity/bijectivity) and quantitative rates in PDE-relevant regimes (furuya2023globally; furuya2024quantitative). Overall, the modern perspective shifts from ”can approximate” to ”how efficiently and under what constraints.” Our work follows this spirit but changes the quantifier structure: the relevant object is not a weight-indexed family across tasks, but the prompt-indexed hypothesis class induced by a fixed executor. Accordingly, the central question becomes whether varying only the prompt yields a family that is dense in $C(K)$, and which structural constraints govern when such switching is possible or impossible.
@@ 2.2 Expressivity and Approximation Results for Transformers
A growing body of theory treats Transformers as explicit function approximators and derives approximation guarantees for attention-based architectures under capacity and regularity assumptions (takakura2023approximation). Subsequent work sharpens these qualitative expressivity statements into quantitative rates, relating approximation error to depth, width, and specific architectural design choices (jiang2024approximation). A closely related direction studies prompt- or context-conditioned function families, where the context acts as an external control signal and expands the realizable set even when the core weights are constrained. In particular, when the context is treated as part of the input, one can connect in-context mappings to universality-style statements for continuous target behaviors (petrov2024prompting; furuya2024transformers).
Complementing approximation guarantees, mechanistic and complexity-oriented analyses link Transformer expressivity to structured computation along the sequence (including intermediate computation traces), and identify task families that expose intrinsic limitations tied to width, depth, and embedding dimension, as well as to parallelizability and finite-precision effects (merrill2023expressive; sanford2023representational; merrill2023parallelism). Our setting is aligned with these perspectives but differs in emphasis: we fix a single executor and vary only the prompt, and we make the switching mechanism explicit through a routing/arithmetic/composition decomposition that is tailored to prompt-controlled computation.
@@ 2.3 Prompting as Adaptation: Soft Prompts, Prefix Tuning, and Prompt Tuning
Prompt-based adaptation is a central parameter-efficient alternative to full fine-tuning, and many approaches can be organized under unified views of updating only a small subset of task-specific parameters (or ”delta” modules) while keeping most weights fixed (ding2023parameter). Within this paradigm, recent work has developed (i) attention-aware soft prompts that are designed to explicitly steer downstream attention patterns (wang2023aprompt), (ii) compositional prompt constructions that reuse shared bases or codebooks to avoid linear parameter growth with prompt length (lin2024accept), and (iii) training objectives that improve robustness and transfer by stabilizing how models rely on prompts across domains (li2024concentrate). Other variants connect prompt tuning to low-rank adaptation, either by introducing low-rank prompt parameterizations that narrow the gap to stronger PEFT baselines while preserving prompt-only deployment (jain2024prompt), or by decomposing prompts into short soft tokens plus auxiliary low-dimensional components (including adaptive, input-dependent forms) to improve efficiency and flexibility (tang2025adept). Analyses of such reparameterizations further suggest that prompt and prefix tuning can encode shared structure that benefits estimation and sample efficiency (le2024revisiting).
In contrast to this primarily algorithmic literature, our focus is expressivity: we treat the prompt as an external parameter indexing a hypothesis class under a fixed executor, and we use a constructive mechanism-level decomposition (routing/arithmetic/composition) to explain how prompt-only switching can be possible in principle, independently of how prompts are found in practice.
@@ 2.4 In-Context Learning and Prompting Theory
A growing literature interprets in-context learning (ICL) as implicit adaptation realized within the forward pass, and asks what learning rule a Transformer effectively implements when provided with demonstrations (akyurek2022learning). Stylized analyses make this view concrete by characterizing task families that are learnable from a finite context window and by relating sample efficiency to the induced algorithmic behavior (garg2022can). Complementary mechanistic and constructive results connect ICL to optimization dynamics, showing that Transformer components can approximate gradient-based updates under suitable parameterizations and that related behaviors can be viewed through meta-optimization lenses (von2023transformers; dai2023can). Other perspectives emphasize representational compression: demonstrations are mapped into a low-dimensional task representation that modulates prediction (hendel2023context), and the optimization view has been extended beyond linear settings via functional gradient descent interpretations (cheng2023transformers). Finally, geometric probes of ICL behavior (e.g., decision-boundary analyses) reveal robustness gaps and irregularities, motivating sharper conditions on what is learned in-context (zhao2024probing).
Our work is aligned with this theoretical agenda in treating prompting as forward-pass computation, but differs in emphasis and framing. We study prompt variation as an external parameter that indexes a hypothesis class under a fixed executor, and we focus on an expressivity (existence) result together with a mechanism-level decomposition (routing/arithmetic/composition), rather than on identifying the specific learning rule induced by demonstrations or on sample-efficiency guarantees for particular task families.
@@ 2.5 Computational and Mechanistic Views: Attention as Retrieval and Transformers as Interpreters
Mechanistic-interpretability research increasingly treats Transformers as decomposable computational systems rather than monolithic black boxes. Automated circuit discovery has supported the view that many attention patterns behave like structured routing mechanisms rather than opaque heuristics (conmy2023towards), and reverse-engineering studies on algorithmic tasks have linked capability shifts to identifiable component-level subroutines (nanda2023progress). End-to-end circuit analyses in language models further attribute substantial portions of model behavior to specific attention-head classes interacting with MLP layers, reinforcing an interpreter-style picture of internal computation (wang2022interpretability).
In parallel, information-flow analyses of factual recall motivate an attention-as-retrieval abstraction, where heads effectively query intermediate representations to extract attributes (geva2023dissecting). Related retrieval framings also appear in long-context modeling, where explicit retrieval mechanisms and self-retrieval formalisms interact with attention as controlled access to stored state (xu2023retrieval; rubin2024retrieval). Finally, causal analyses of induction heads in ICL provide fine-grained evidence that particular heads implement pattern-matching circuits, strengthening a routing-centric interpretation of in-context behavior (crosbie2025induction).
Our lemma-based decomposition can be viewed as a proof-level abstraction of these mechanistic roles: prompt slots serve as addressable memory, attention implements selective routing/retrieval, the FFN implements local computation on retrieved fragments, and depth-wise composition accumulates these operations into a multi-step program execution.
@@ 3.3 The Prompt-UAT objective
We allow $\mathcal{P}=\mathbb{R}^{L_{\max}\times d_{p}}$ without norm constraints to establish an existential expressivity result; quantitative trade-offs (prompt norm/precision versus $\varepsilon$) are deferred.
@@ 4.1 Architecture
A prompt is a length-$L$ token sequence or matrix) $P=[p_{1};\ldots;p_{L}]\in\mathbb{R}^{L\times D}$ with $p_{i}\in\mathbb{R}^{D}$. An input $x\in\mathcal{X}$ is embedded into a single input token $u(x)=E_{x}x+b_{x}\in\mathbb{R}^{D}$, where $E_{x}\in\mathbb{R}^{D\times d}$ and $b_{x}\in\mathbb{R}^{D}$. We also introduce an initial output token $o\in\mathbb{R}^{D}$. Letting the sequence length be $n=L+3$, the initial state is
@@
$\mathcal{A}_{\tau}(Z):=\mathcal{S}_{\tau}\!\Big(\frac{(ZW_{Q})(ZW_{K})^{\top}}{\sqrt{D}}\Big)\,(ZW_{V}).$
(Equivalently, the $\sqrt{D}$ scaling can be absorbed into $\tau$; we write it explicitly for clarity.) Unless stated otherwise, we use full self-attention; in a causal variant, the output token can still attend to the prompt prefix.
@@ 4.2 Viewing the prompt as key–value memory
$\mathcal{R}_{\tau}(q;P):=\mathcal{S}_{\tau}(qK_{P}^{\top})\,V_{P}\in\mathbb{R}^{D}.$
Then $\mathcal{R}_{\tau}(q;P)$ is a convex combination of rows of $V_{P}$. Hence, if a sufficient score margin is ensured, a small $\tau$ yields near one-hot routing that concentrates on a single slot (Routing Lemma).
@@ 7 Construction: Prompt-Encoded MLP Execution
This section presents only the skeleton of the interpreter (executor) construction. The key idea is to first approximate the target function $g$ by a ReLU MLP $N_{g}$, and then design a fixed Transformer $\theta^{\ast}$ that uniformly emulates the forward pass of $N_{g}$ by reading $N_{g}$’s parameters stored in the prompt. Throughout this section, the input domain is $\mathcal{X}\subset\mathbb{R}^{d}$ (compact) as in Section 3, and we fix a maximum prompt length $L_{\max}$ (hence the available memory). Different tasks may use different numbers of slots, implemented via zero-padding within length $L_{\max}$.
@@ 7.1 Encoding: Prompt as Parameter Memory
Let $N:\mathbb{R}^{d}\to\mathbb{R}$ be a depth-$S$ ReLU MLP defined by
@@
$h^{(\ell)}(x)$ $=\sigma\!\big(W^{(\ell)}h^{(\ell-1)}(x)+b^{(\ell)}\big),\quad\ell=1,\dots,S-1,$
$N(x)$ $=a^{\top}h^{(S-1)}(x)+c,$
where $\sigma=\mathrm{ReLU}$. We assume that for some $\Lambda>0$ all parameters are uniformly bounded in the entrywise max norm: $\|W^{(\ell)}\|_{\max}\leq\Lambda$, $\|b^{(\ell)}\|_{\max}\leq\Lambda$, $\|a\|_{\max}\leq\Lambda$, and $|c|\leq\Lambda$.
@@
Consider a depth-$S=3$ ReLU MLP $N(x)=a^{\top}\sigma(Wx+b)+c$. A concrete choice of chunking is to store (i) each row $W_{r:}$ and the corresponding bias entry $b_{r}$ as one fragment, (ii) the output weights $a_{r}$ as one fragment per hidden unit, and (iii) the scalar $c$ as a final fragment. Thus $J$ can be indexed by hidden-unit IDs together with a small tag indicating whether the payload is $(W_{r:},b_{r})$ or $a_{r}$, plus one slot for $c$; the addressing map $\mathrm{addr}:J\to[L]$ places these fragments into predetermined prompt slots. During execution, the work token retrieves the fragments for unit $r$, computes $\langle W_{r:},x\rangle+b_{r}$, applies ReLU, multiplies by $a_{r}$, and accumulates across $r$; repeating this logic realizes the layerwise macro-step in Definition 7.7.
@@
Fix a finite index set $\mathcal{J}$ and a chunking map $j\mapsto\mathrm{payload}_{j}(N)\in\mathbb{R}^{q}$ that extracts parameter fragments of $N$ (e.g., blocks of weights/biases or individual rows/columns). Fix an addressing function $\mathrm{addr}:\mathcal{J}\to[L]$ that assigns each fragment index $j$ to a prompt slot.
@@
(after padding to $D\times D$). Then for any prescribed keys $\kappa(i)\in\mathbb{R}^{d_{k}}$ and any prescribed values $\nu(y_{i})\in\mathbb{R}^{d_{v}}$, the choice
$p_{i}=\begin{bmatrix}\kappa(i)\\ \nu(y_{i})\\ 0\end{bmatrix}\in\mathbb{R}^{D}$
satisfies $(p_{i}W_{K})=\kappa(i)$ and $(p_{i}W_{V})=\nu(y_{i})$ simultaneously.
@@
Fix an addressing (key) scheme $\kappa:[L]\to\mathbb{R}^{d_{k}}$ and a value embedding map $\nu:\mathbb{R}^{q}\to\mathbb{R}^{d_{v}}$. Assuming Lemma 7.3, define the encoding $p_{N}=\mathrm{Enc}(N)\in\mathbb{R}^{L\times D}$ as follows: for each $j\in\mathcal{J}$, let $i=\mathrm{addr}(j)$ and set the value stored at slot $i$ to be $\nu(\mathrm{payload}_{j}(N))$, while the slot key is fixed as $\kappa(i)$.

### v2

@@ 2.1 Universal Approximation and Expressivity Foundations
Universal approximation characterizes expressivity via density of realizable functions in spaces such as $C(K)$ for compact $K$. Recent theory refines the classical UAT by providing quantitative rates tied to architectural capacity and function-space regularity (e.g., Besov-type smoothness for ReLU networks) (Siegel, 2023), and by identifying minimal width conditions and sharp thresholds under which universality holds or fails even on compact domains (Kim et al., 2023). Extensions to broader settings consider uniform approximation beyond compact domains, where tail control becomes essential (Bournez et al., 2025), and alternative parameterizations that preserve approximation goals while altering the mechanism (Liu et al., 2024). Related ideas appear in operator learning, which elevates universality to maps between function spaces and connects approximation theory to PDE solution operators (Kovachki et al., 2023; Furuya et al., 2023, 2024b). Our work follows this modern viewpoint but changes the quantifier structure: rather than a weight-indexed family across tasks, we study the prompt-indexed hypothesis class induced by a fixed executor, asking when varying only the prompt yields density in $C(K)$ and what structural constraints govern such switching.
@@ 2.2 Expressivity and Approximation Results for Transformers
A growing theory treats Transformers as explicit function approximators and derives approximation guarantees under capacity and regularity assumptions (Takakura and Suzuki, 2023; Jiang and Li, 2024). Closely related are results on prompt- or context-conditioned families, where context acts as a control signal that expands the realizable set even with constrained core weights (Petrov et al., 2024; Furuya et al., 2024a). Complementary complexity- and mechanism-oriented analyses link Transformer expressivity to structured computation along the sequence and identify intrinsic limitations tied to width, depth, embedding dimension, and finite-precision or parallelizability constraints (Merrill and Sabharwal, 2023a; Sanford et al., 2023; Merrill and Sabharwal, 2023b). Our setting aligns with these perspectives but focuses on a single fixed executor with prompt-controlled computation, made explicit via a routing/arithmetic/composition decomposition.
@@ 2.3 Prompting as Adaptation: Soft Prompts, Prefix Tuning, and Prompt Tuning
Prompt-based adaptation is a widely used parameter-efficient alternative to full fine-tuning, and many methods can be viewed as updating a small task-specific parameterization while keeping most weights fixed (Ding et al., 2023). Recent variants design prompts to steer attention patterns (Wang et al., 2023), construct compositional or codebook-based prompts to control parameter growth (Lin et al., 2024), and stabilize prompt reliance to improve robustness and transfer (Li et al., 2024). Other work connects prompt tuning to low-rank reparameterizations and hybrid prompt modules that improve efficiency and flexibility (Jain et al., 2024; Tang et al., 2025), with analyses suggesting shared structure that benefits estimation and sample efficiency (Le et al., 2024). In contrast, we focus on expressivity: prompts index a hypothesis class under fixed weights, and we provide a constructive mechanism-level account of when prompt-only switching is possible in principle.
@@ 2.4 In-Context Learning and Prompting Theory
Theoretical work on in-context learning (ICL) interprets prompting as forward-pass adaptation and asks what algorithmic behavior is implemented when demonstrations are provided (Akyürek et al., 2022; Garg et al., 2022). Constructive and mechanistic analyses connect ICL to optimization dynamics and meta-learning views, showing that Transformer components can approximate gradient-based updates under suitable parameterizations (Von Oswald et al., 2023; Dai et al., 2023), and other perspectives emphasize compression into low-dimensional task representations (Hendel et al., 2023) or extend optimization views beyond linear settings (Cheng et al., 2023). Empirical and geometric probes further highlight robustness gaps and irregularities, motivating sharper conditions on what is learned in-context (Zhao et al., 2024). We are aligned with this agenda in treating prompting as computation, but we study an expressivity (existence) result for a fixed executor and make the switching mechanism explicit through routing/arithmetic/composition, rather than characterizing a particular learning rule or sample-efficiency regime.
@@ 2.5 Computational and Mechanistic Views: Attention as Retrieval and Transformers as Interpreters
Mechanistic-interpretability work increasingly models Transformers as decomposable computational systems, where attention behaves like structured routing and circuits implement recognizable subroutines on algorithmic tasks (Conmy et al., 2023; Nanda et al., 2023; Wang et al., 2022). In parallel, attention-as-retrieval abstractions connect factual recall and long-context behavior to controlled access to stored state or self-retrieval formalisms (Geva et al., 2023; Xu et al., 2023; Rubin and Berant, 2024), and causal analyses of induction heads provide fine-grained evidence for routing-centric interpretations of ICL (Crosbie and Shutova, 2025). Our lemma-based decomposition can be seen as a proof-level abstraction of these roles: prompt slots act as addressable memory, attention implements selective retrieval, the FFN performs local computation on retrieved fragments, and depth-wise composition accumulates these operations into a multi-step program.
@@ 3.3 The Prompt-UAT objective
We allow the ambient prompt space $\mathcal{P}=\mathbb{R}^{L_{\max}\times d_{p}}$ without norm constraints to state a purely existential expressivity result; quantitative trade-offs (prompt norm/precision versus $\varepsilon$) are deferred. In the actual executor-level construction and error analysis, however, we work with an admissible subset $\mathcal{P}_{\mathrm{adm}}\subset\mathcal{P}$ produced by the encoding scheme, on which all stored records are uniformly bounded.
@@ 4.1 Architecture
A prompt is a length-$L$ token sequence (matrix) $P=[p_{1};\ldots;p_{L}]\in\mathbb{R}^{L\times D}$ with $p_{i}\in\mathbb{R}^{D}$. An input $x\in\mathcal{X}$ is embedded into a single input token $u(x)=E_{x}x+b_{x}\in\mathbb{R}^{D}$, where $E_{x}\in\mathbb{R}^{D\times d}$ and $b_{x}\in\mathbb{R}^{D}$. We also introduce an initial output token $o\in\mathbb{R}^{D}$ and a work token $w\in\mathbb{R}^{D}$ (initialized to $w=\mathbf{0}$), which serves as a scratch register for intermediate computations. Letting the sequence length be $n=L+3$, the initial state is
@@
$\mathcal{A}_{\tau}(Z)$ $=\mathcal{S}_{1}\!\Bigg(\frac{(ZW_{Q})(ZW_{K})^{\top}}{\tau\sqrt{D}}\Bigg)\,(ZW_{V})$
$=\mathcal{S}_{\tau}\!\Bigg(\frac{(ZW_{Q})(ZW_{K})^{\top}}{\sqrt{D}}\Bigg)\,(ZW_{V}).$
Unless stated otherwise, we use full self-attention; in a causal variant, the output token can still attend to the prompt prefix.
@@ 4.2 Viewing the prompt as key–value memory
$\mathcal{R}_{\tau}(q;P):=\mathcal{S}_{\tau}\!\left(\frac{qK_{P}^{\top}}{\sqrt{D}}\right)\,V_{P}\in\mathbb{R}^{D}.$
Since $\mathcal{S}_{\tau}(\cdot)$ outputs a probability vector (nonnegative entries summing to $1$), $\mathcal{R}_{\tau}(q;P)$ is a convex combination of the rows of $V_{P}$. Hence, if a sufficient score margin is ensured, taking $\tau$ small yields near one-hot routing that concentrates on a single slot (Routing Lemma).
@@ 7 Construction: Prompt-Encoded MLP Execution
This section presents the skeleton of the interpreter (executor) construction. The key idea is to first approximate a target function $g$ by a ReLU MLP $N_{g}$, and then design a fixed Transformer $\theta^{\ast}$ that uniformly emulates the forward pass of $N_{g}$ by reading $N_{g}$’s parameters stored in the prompt. Throughout this section, the input domain is $\mathcal{X}\subset\mathbb{R}^{d}$ (compact) as in Section 3. We fix a maximum prompt length $L_{\max}$ (hence the available memory) and represent every prompt as a matrix in $\mathbb{R}^{L_{\max}\times D}$, using zero-padding in unused slots. Let $L\leq L_{\max}$ denote the number of nonzero (used) slots.
@@ 7.1 Encoding: Prompt as Parameter Memory
Let $N:\mathbb{R}^{d}\to\mathbb{R}$ be a depth-$S$ ReLU MLP, where (by convention) $S$ counts the input layer, the (S2)(S\!-\!2) hidden ReLU layers, and the output layer. Let $m_{\ell}$ denote the width of hidden layer $\ell$ and set $m_{0}=d$. Define
@@
$h^{(\ell)}(x)$ $=\sigma\!\big(W^{(\ell)}h^{(\ell-1)}(x)+b^{(\ell)}\big),\quad\ell=1,\dots,S-2,$
$N(x)$ $=W^{(S-1)}h^{(S-2)}(x)+b^{(S-1)},$
where $\sigma=\mathrm{ReLU}$, $W^{(\ell)}\in\mathbb{R}^{m_{\ell}\times m_{\ell-1}}$ and $b^{(\ell)}\in\mathbb{R}^{m_{\ell}}$ for $\ell\leq S-2$, and the output layer satisfies $W^{(S-1)}\in\mathbb{R}^{1\times m_{S-2}}$, $b^{(S-1)}\in\mathbb{R}$. We assume that for some $\Lambda>0$ all parameters are uniformly bounded in the entrywise max norm: $\|W^{(\ell)}\|_{\max}\leq\Lambda$, $\|b^{(\ell)}\|_{\max}\leq\Lambda$ for all $\ell\in\{1,\dots,S-1\}$.
@@
Consider a depth-$S=3$ ReLU MLP with one hidden layer:
$N(x)=a^{\top}\sigma(Wx+b)+c,$
which matches Definition 7.1 by identifying $W^{(1)}=W$, $b^{(1)}=b$, $W^{(2)}=a^{\top}$, and $b^{(2)}=c$. A concrete choice of chunking is to store (i) each row $W_{r:}$ and the corresponding bias entry $b_{r}$ as one fragment, (ii) the output weight $a_{r}$ as one fragment per hidden unit, and (iii) the scalar $c$ as a final fragment. Thus $\mathcal{J}$ can be indexed by hidden-unit IDs together with a small tag indicating whether the payload is $(W_{r:},b_{r})$ or $a_{r}$, plus one slot for $c$; the addressing map $\mathrm{addr}:\mathcal{J}\to[L_{\max}]$ places these fragments into predetermined prompt slots. During execution, the work token retrieves the fragments for unit $r$, computes $\langle W_{r:},x\rangle+b_{r}$, applies ReLU, multiplies by $a_{r}$, and accumulates across $r$; repeating this logic realizes the layerwise macro-step in Definition 7.7.
@@
Fix a finite index set $\mathcal{J}$ and a chunking map $j\mapsto\mathrm{payload}_{j}(N)\in\mathbb{R}^{q}$ that extracts parameter fragments of $N$ (e.g., blocks of weights/biases or individual rows/columns). Fix an addressing function $\mathrm{addr}:\mathcal{J}\to[L_{\max}]$ that assigns each fragment index $j$ to a prompt slot.
@@
(after padding to $D\times D$). Let each prompt token be a row vector $p_{i}\in\mathbb{R}^{1\times D}$ and define
$p_{i}=\big[\ \kappa(i)^{\top}\ \ \nu(y_{i})^{\top}\ \ 0^{\top}\ \big]\in\mathbb{R}^{1\times D},$
where $\kappa(i)\in\mathbb{R}^{d_{k}}$ and $\nu(y_{i})\in\mathbb{R}^{d_{v}}$ are prescribed. Then $(p_{i}W_{K})_{1:d_{k}}=\kappa(i)$ and $(p_{i}W_{V})_{(d_{k}+1):(d_{k}+d_{v})}=\nu(y_{i})$ simultaneously.
@@
Fix an addressing (key) scheme $\kappa:[L_{\max}]\to\mathbb{R}^{d_{k}}$ and a value embedding map $\nu:\mathbb{R}^{q}\to\mathbb{R}^{d_{v}}$. Assuming Lemma 7.3, define the encoding $p_{N}=\mathrm{Enc}(N)\in\mathbb{R}^{L_{\max}\times D}$ as follows: for each $j\in\mathcal{J}$, let $i=\mathrm{addr}(j)$ and set the value stored at slot $i$ to be $\nu(\mathrm{payload}_{j}(N))$, while the slot key is fixed as $\kappa(i)$.